### v2

@@ A.5 Tight Example on Demand-Approximate CE
The prices of type-O, P, and S goods are 00. and prices of type-R goods are $\frac{n+1}{n}$. In this case, the type-A agent will get the type-O good and type-P goods, resulting in a utility of $n+1$. Each type-B agent receives $\frac{n}{n+1}$ portion of the corresponding type-R good, leading to a utility of $\frac{n}{n+1}$. The type-C agent acquires $1$ portion of the corresponding type-S good and the rest of type-R goods, resulting in a utility of $1+\frac{n}{n+1}$.
@@
A.8 Tight Example of Theorem 1.3
The following example shows the bound in Theorem 1.3 is tight.
Example A.9.
Consider the case that there are $n$ agents and $n$ goods. All agents have bounded linear utilities. Additionally, all agents have the same preferences ($v_{ij}=1$) on all goods except for the first good. For the first good, only the first agent likes the first good ($v_{11}=0$) and all other agents don’t like this good $v_{i1}=0$ for $i\neq 1$. Additionally, we assume the upper bound on the utility of each good of each agent is $1$, $l_{ij}=1$ for all $i$ and $j$.
In this case, for the maximal Nash welfare solution, the utilities of all agents equal $1$. However, there exists a competitive equilibrium, at which the price of the first good equals 00 and the prices for all other goods equal $\frac{n}{n-1}$. Then, the first agent can get the first good and $\frac{n-1}{n}$ fraction of the second good, and achieve a utility of $1+\frac{n-1}{n}$ which tends to $2$ in the case $n\rightarrow\infty$.
A.9 Example with Leontief-free utility functions
@@
Consider the scenario with $n$ goods and $n+1$ agents. Agent $i=1,2,\cdots,n$ only likes good $i$, has budget $b_{i}=1$ and a linear utility function $u_{i}(x_{i})=x_{ii}$. Agent $n+1$ has a substantially larger budget $b_{k+1}=B$, and a utility function $\min\{\sum_{i=1}^{k}x_{(n+1)i},(n-1)(1-\epsilon)\}$ with a small $\epsilon>0$.
@@
There is a competitive equilibrium where the prices are $p_{1}=1$ and $p_{2}=p_{3}=\cdot=p_{n}=\frac{1}{\epsilon}$. Agent $1$ obtains $x_{11}=1$, the other agents get $x_{ii}=\epsilon$ for $i=2,\cdots,n$, and agent $n+1$ gets $x_{(n+1)1}=0$ and $x_{(n+1)j}=1-\epsilon$ for $j=2,\cdots,n$. The competitive equilibrium holds if $B>\frac{(n-1)(1-\epsilon)}{\epsilon}$.
@@
On the other hand, in the Nash welfare maximizing allocation, when $B$ is big enough, agent $n+1$ receives (approximately) $\frac{(1-\epsilon)(n-1)}{n}$ of each good, with the remaining goods shared among the remaining agents. Agent $1$ receives a utility of $\epsilon+\frac{1-\epsilon}{n}$, which is significantly less than the utility received in the competitive equilibrium.
@@
In this case, agent $n+1$’s utility doesn’t satisfy property ($\Sigma$). When $\mathbf{p}=\mathbf{0}$, $x_{n+1}=(0,1-\epsilon,1-\epsilon,\cdots,1-\epsilon)\in\mathcal{GD}_{n+1}(\mathbf{p},B)$. However, when $\mathbf{p}=(1,\frac{1}{\epsilon},\frac{1}{\epsilon},\cdots,\frac{1}{\epsilon})$, for any $x_{n+1}\in\mathcal{GD}_{n+1}(\mathbf{p},B)$, $x_{(n+1)1}>0$, which violates property ($\Sigma$).
A.10 Example with matching utility functions
The following example illustrates that matching utilities are not Gale-substitutes in general.
Example A.11.
@@ A.10 Example with matching utility functions
Consider the scenario with $3$ goods and an agent with unit budget. The agent’s utility is a matching utilities, with preferences for the goods being $1$, $2.5$, and $0.7$ respectively. The initial prices for the goods are $0.3$, $1.2$, and $0.1$. In this case, the Gale demand is
${\arg\max}_{x=(x_{1},x_{2},x_{3}):x_{1}+x_{2}+x_{3}\leq 1}\{\log(x_{1}+2.5x_{2}+0.7x_{3})-(0.3x_{1}+1.2x_{2}+0.1x_{3})\},$
@@
which gives the solution $(0,103/198,95/198)$. Now we keep the prices of the first two goods unchanged and increase the price of the third good from $0.1$ to $0.6$. By calculation, under the new prices, the new Gale demand is $(5/9,4/9,0)$. The Gale demand of good $2$ decreases, which implies that it is not substitutes.
Appendix B Proof of Lemma 8.1
See 8.1
Proof.
First, assume that $\pi\in\mathbb{R}_{\geq 0}^{\ell}$ is an optimal solution to $\operatorname{Dual}(x)$ for a given $x\in\mathbb{R}^{m}$. We need to show that $g=B^{\top}\pi$ is a supergradient at $x$, that is, $\varphi(t)\leq\varphi(x)+\left\langle g,t-x\right\rangle$ for any $t\in\mathcal{D}$.
The programs $\operatorname{Dual}(x)$ and $\operatorname{Dual}(t)$ have the same feasible region, with linear objective functions $\mathbf{B}x+b$ and $\mathbf{B}t+b$, respectively. By weak duality, the cost of every feasible dual solution is at least the primal optimum value, thus, $\varphi(t)\leq t^{\top}\mathbf{B}\pi+b^{\top}\pi=g^{\top}t+b^{\top}\pi$. The statement follows as
$\varphi(t)\leq g^{\top}t+b^{\top}\pi=g^{\top}(t-x)+x^{\top}\mathbf{B}\pi+b^{\top}\pi=\varphi(x)+g^{\top}(t-x)\,,$
noting also that $\varphi(x)=x^{\top}\mathbf{B}\pi+b^{\top}\pi$ since $\pi$ is an optimal solution to $\operatorname{Dual}(x)$, and $\varphi(x)$ is the optimum value.
Conversely, we need to show that if $g\in\partial^{\wedge}\varphi(x)$, then there exists a dual optimal solution $\pi$ such that $g=\mathbf{B}^{\top}\pi$. Let $z$ be a primal optimal solution to $\operatorname{Primal}(x)$. Using complementary slackness, we can turn $\operatorname{Dual}(x)$ into a feasibility system. Namely, $g=\mathbf{B}^{\top}\pi$ for a dual optimal solution if and only if
$\mathbf{B}^{\top}\pi$ $=g$ (12)
$\mathbf{A}_{i}^{\top}\pi$ $=c_{i}\,,\quad\forall i:z_{i}>0$
$\mathbf{A}_{i}^{\top}\pi$ $\geq c_{i}\,,\quad\forall i:z_{i}=0$
$\pi_{j}$ $=0\,,\quad\forall j:(\mathbf{A}z)_{j}<(\mathbf{B}x+b)_{j}$
$\pi_{j}$ $\geq 0,\quad\forall j:(\mathbf{A}z)_{j}=(\mathbf{B}x+b)_{j}\,.$
Here, $\mathbf{A}_{i}$ denotes the $i$-th column of $\mathbf{A}$. By Farkas’s lemma, if (12) is not feasible, then the following system has a feasible solution:
$(\mathbf{A}w)_{j}+(\mathbf{B}r)_{j}$ $\leq 0\,,\quad\forall j:(\mathbf{A}z)_{j}=(\mathbf{B}x+b)_{j}$
$w_{i}$ $\geq 0\,,\quad\forall i:z_{i}=0$
$c^{\top}w+g^{\top}r$ $>0\,.$
For a small enough $\varepsilon>0$, $z+\varepsilon w$ is feasible to the system $\operatorname{Primal}(x-\varepsilon r)$. Hence, for $t=x-\varepsilon r$, $\varphi(t)\geq c^{\top}(z+\varepsilon w)=c^{\top}z+\varepsilon c^{\top}w$. Using $c^{\top}z=\varphi(x)$ by the choice of $z$ as an optimal solution to $\operatorname{Primal}(x)$, and the bound $c^{\top}w>-g^{\top}r$, we get $\varphi(t)>\varphi(x)-\varepsilon g^{\top}r=\varphi(x)+\left\langle g,t-x\right\rangle$, a contradiction to $g\in\partial^{\wedge}\varphi(x)$. ∎

### v3

@@ E.2 Non-monotonicity of Gale Demands
At the thrifty competitive equilibrium, the prices of type-O, P, and S goods are 0, while the prices of type-R goods are $\frac{n+1}{n}$. In this case, the type-A agent will get the type-O good and type-P goods, resulting in a utility of $n+1$. Each type-B agent receives $\frac{n}{n+1}$ portion of the corresponding type-R good, leading to a utility of $\frac{n}{n+1}$. The type-C agent acquires $1$ portion of the corresponding type-S good and the rest of type-R goods, resulting in a utility of $1+\frac{n}{n+1}$.
@@
E.5 Example with Leontief-free utility functions
@@
Consider the scenario with $n$ goods and $n+1$ agents. Agent $i=1,2,\cdots,n$ only likes good $i$, has budget $b_{i}=1$ and a linear utility function $u_{i}(\mathbf{x}_{i})=x_{ii}$. Agent $n+1$ has a substantially larger budget $b_{n+1}=B$, and a utility function $\min\{\sum_{j=1}^{n}x_{(n+1)j},(n-1)(1-\varepsilon)\}$ with a small $\varepsilon>0$.
There is a competitive equilibrium where the prices are
$p_{1}=1\text{ and }p_{2}=p_{3}=\cdot=p_{n}=\frac{1}{\varepsilon}\,.$
@@
Agent $1$ obtains $x_{11}=1$, the other agents get $x_{ii}=\varepsilon$ for $i=2,\cdots,n$, and agent $n+1$ gets $x_{(n+1)1}=0$ and $x_{(n+1)j}=1-\varepsilon$ for $j=2,\cdots,n$. The competitive equilibrium holds if $B>\frac{(n-1)(1-\varepsilon)}{\varepsilon}$.
@@
On the other hand, in the Nash welfare maximizing allocation, when $B$ is big enough, agent $n+1$ receives (approximately) $\frac{(1-\varepsilon)(n-1)}{n}$ of each good, with the remaining goods shared among the remaining agents. Agent $1$ receives a utility of $\varepsilon+\frac{1-\varepsilon}{n}$, which is significantly less than the utility received in the competitive equilibrium.
@@
In this case, agent $n+1$’s utility doesn’t satisfy property ($\Sigma$). When $\mathbf{p}=\mathbf{0}$,
$\mathbf{x}_{n+1}=(0,1-\varepsilon,1-\varepsilon,\cdots,1-\varepsilon)\in\mathcal{GD}_{n+1}(\mathbf{p},B)\,.$
However, when $\mathbf{p}=(1,\frac{1}{\varepsilon},\frac{1}{\varepsilon},\cdots,\frac{1}{\varepsilon})$,
$\text{for any }\mathbf{x}_{n+1}\in\mathcal{GD}_{n+1}(\mathbf{p},B),\ \ x_{(n+1)1}>0,$
which violates property ($\Sigma$).
Observe that we can make the budgets of all agents the same by creating $B$ copies of agent $n+1$, each with the utility function $\min\{\sum_{j=1}^{n}x_{(n+1)j},(n-1)(1-\varepsilon)/B\}$.
E.6 Example with matching utility functions
The following example shows that matching utility functions (i.e., those with a unit-demand constraint) are not Gale-substitutes in general. This also implies that utility functions with more general constraints (as in Garg et al. (2022b), Jalota et al. (2023)) need not satisfy the Gale-substitutes property either.
Example E.7.
@@ E.6 Example with matching utility functions
Consider a market with $3$ goods and an agent with unit budget. The agent has a matching utility function (i.e., a unit-demand constraint), with valuations of $1$, $2.5$, and $0.7$ for goods 1, 2, and 3, respectively. The initial prices vector is $(0.3,1.2,0.1)$. In this case, the Gale demand is
${\arg\max}_{\mathbf{x}:x_{1}+x_{2}+x_{3}\leq 1}\{\log(x_{1}+2.5x_{2}+0.7x_{3})-(0.3x_{1}+1.2x_{2}+0.1x_{3})\},$
@@
which yields the solution $(0,103/198,95/198)$. Now, keeping the prices of the first two goods unchanged, we increase the price of the third good from $0.1$ to $0.6$. Under these new prices, the Gale demand becomes $(5/9,4/9,0)$. Since the demand for good $2$ decreases as the price of good $3$ increases, this shows that the utility function does not satisfy the Gale-substitute property.